The production function of producing computer software diskettes (Q) is given as: q=LK^1/2. The input K is fixed at 100. The price of K is $2 and the price of L is $50. The short run total cost function is: TC= ___+___*Q, where Q is the Level of ouput.
Answer #1

Solution:

Here the data given is:

  • K = 100
  • Wage rate = $50
  • Capital rate = $2
  • Q = LK^1/2

Now, q = LK^1/2

Q = L * (100)^1/2

Q = L * 10

L = Q/10

Total cost = fixed cost (FC) + variable cost(VC)

TC = K*capital rate + wage rate*L

TC = 100*2 + 50* Q/10

TC = 200 + 5Q

Hence , short run total cost function ( TC ) = 200 + 5Q